OpenStudy (anonymous):

Suppose f(x) is continuous on [2,8] and −4≤f′(x)≤5 for all x in (2,8). Use the Mean Value Theorem to estimate f(8)−f(2).

OpenStudy (anonymous):

by mvt \[\frac{f(8)-f(2)}{8-2}=f'(c)\] for some c in (2,8)

OpenStudy (anonymous):

so you have \[-4\leq\frac{f(8)-f(2)}{6}\leq 5\]

OpenStudy (anonymous):

therefore \[-24\leq f(8)-f(2)\leq 30\]